OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -10 - 2 + 6 + 14 + ... + 110

OpenStudy (anonymous):

as -2 -(-10) =8 6- (-2) =8 thus the above pattern is in AP

OpenStudy (anonymous):

AP means Arithmetic sequence hence here the first term is -10 and the common difference is +8 first we have to find the number of terms involved say n

OpenStudy (anonymous):

how would I find that?

OpenStudy (anonymous):

here an the nth tern is 110 so we have 110 = -10 + (n-1)*8 thus n-1 =(110+10)/8 or n-1 =15 hence n =16 that is to saty 110 is the 16th term of the given pattern

OpenStudy (anonymous):

so far I have \[\sum_{n}^{n=0} -10 + 8 (n)\]

OpenStudy (anonymous):

so the summation all the numbers in the given pattern is yup correct the limits are now from 0 to 15 (instead of 16 as the start point is 0) as we know the value of n

OpenStudy (anonymous):

i hope you can do futher this may result as -10*16 + 8*(1+2+3+4...........15) =-160 + 8*15*16/2 and so on ...

OpenStudy (anonymous):

\[\sum_{n=0}^{15} -10 + 8n or \sum_{n=0}^{15} -80n\]

OpenStudy (anonymous):

first one..

OpenStudy (anonymous):

okay so you don't add them together, just leave them as they are?

OpenStudy (anonymous):

as they are not alike terms..

OpenStudy (anonymous):

now the sum equals -10*16 + 8*(1+2+3+4...........15) =-160 + 8*15*16/2 and so on ...

OpenStudy (anonymous):

i hope that you will be able to carry on further...

OpenStudy (anonymous):

I understand now. Sorry for being difficult I forget the simple things sometimes :/ oops

OpenStudy (nirmalnema):

this is an AP in which first term is -10 and last term is 110 having common difference = -2-(-10)=8 first u have to find total no of terms any term=a+(n-1)d n=no of term a=first term d=common difference 110=-10+(n-1)8 n=16 sum of all terms= n/2 * (2a+(n-1)d) 16/2 * (-20+(15)(8)) 8*100 =800

OpenStudy (anonymous):

@yesy7 its ok

OpenStudy (anonymous):

Thank you @nirmalnema for explaining as well :)

OpenStudy (nirmalnema):

its ok..!!

OpenStudy (anonymous):

@yesy7 what nirmelnema has done is also an alternate way of finding the required sum further since you know the last term we can alternatively (rather more easy) use sum =n/2 *(first term +last term) hence sum =16/2 (-10 +110) = 8* 100 =800

OpenStudy (anonymous):

Okay, cool!